IF a current in a bulb is 3A . if it has 40 ohm resistance then what is the power generated by it.
 

Grade:11

1 Answers

Ursula Arthur
2 years ago
Current is 3A and resistance is 40 ohms from power(P)=VI ...(equation 1)
But V=I R .....(equation 2)
We substitute equation 2 into equation 1.
P=I (I R)
P=3(3×40)
P=3×120
P=360 Watt